The base of a triangle is 8 cm greater than the height. The area is 42 cm^2. Find the height and the length of the base. I need help with this please. Thank you.
Answer by mananth(16949)

let thge height be x
the base = x+8
Area of trangle = 1/2 * base * height
1/2 * x*(x+8) = 42
1/2 (x^2+8x)=42
multiply equation by 2
x^2+8x=42*2
x^2+8x=84
x^2+8x-84=0
x^2+14x-6x-84=0
x(x+14)-6(x+14)=0
(x-6)(x+14)=0
x=6 OR x= -14
x=6 is the height
base = x+8 = 6+8 =14